OpenStudy (anonymous):

if 18,000 divided by n =300 then what is the value of n ?

OpenStudy (anonymous):

helppp plase

OpenStudy (anonymous):

\[\frac{ 18,000 }{ x }=300\]

OpenStudy (anonymous):

Like that?

OpenStudy (anonymous):

18000/x=300 x=18000/300 x=60